# Massive Higher-Spin Fields in the Fractional Quantum Hall Effect

### v1

@@ Appendix A Proca-like actions for massive particles of any spin
The value of $c_{2}$ fixed by Eq. (35) is precisely that setting to zero the right-hand side. This allows one to conclude $\partial_{\mu}\partial\cdot B=0$ and, eventually, $\partial\cdot B=0$ up to constant functions that do not affect the propagated local degrees of freedom. This on-shell constraint is crucial because, after the gauge fixing setting the scalar to zero, the spin-$2$ gauge parameter satisfies $\partial\cdot\xi=0$, so that the divergence of the vector potential becomes gauge invariant. On the other hand, the latter is forced to vanish by the equations of motion, so that one can gauge away the remaining portion of the vector potential by using the algebraic gauge transformation generated by $\xi_{\mu}$. After one reaches the condition $B_{\mu}=0$ by combining the equations of motion and a gauge fixing, Eq. (38) reduces to the Fierz-Pauli equation (10).
Appendix B Non-relativistic limit
In this section we look at a particular Galilean limit of the higher-spin actions presented in the previous section. We consider the examples of fields of spin two and four and restrict the analysis to $D=2+1$ dimensions. The relativistic indices will be split in the form $\mu=\{0,a\}$ and we consider coordinates $x^{\mu}=\{x^{0}\equiv t,x^{a}\}$ so that the Minkowski metric and its inverse are given by $\eta_{\mu\nu}={\rm diag}(-c^{2},1,1)$ and $\eta^{\mu\nu}={\rm diag}(-1/c^{2},1,1)$, respectively. Let us consider now an explicit traceless parametrisation of the fields entering the action (34) up to spin-$4$, which can be written in terms of a set of six completely symmetric and traceless tensors $\{S_{abcd},S_{abc},S_{ab},v_{abc},v_{ab},\phi_{ab}\}$, four vectors $\{S_{a},v_{a},\phi_{a},\alpha_{a}\}$ and five scalar fields $\{\beta,\alpha,v,\phi,S\}$ as follows. In order to avoid confusion, we rename the fields as
$\varphi_{\mu(0)}=S,\quad\varphi_{\mu(1)}=B_{\mu},\quad\varphi_{\mu(2)}=h_{\mu\nu},\quad\varphi_{\mu(3)}=\chi_{\mu\nu\rho},\quad\varphi_{\mu(4)}=\Phi_{\mu\nu\rho\sigma},\quad$ (41)
For the spin-$1$ and spin-$2$ auxiliary fields we define
$B_{\mu}=\begin{pmatrix}B_{0}\\ B_{a}\end{pmatrix}\equiv\begin{pmatrix}v\\ S_{a}\end{pmatrix},\qquad h_{\mu\nu}=\left(\begin{array}[]{c|c}h_{00}&h_{0b}\\ \hline\cr h_{a0}&h_{ab}\end{array}\right)\equiv\left(\begin{array}[]{c|c}\phi&v_{b}\\ \hline\cr v_{a}&\frac{1}{2c^{2}}\phi\,\delta_{ab}+S_{ab}\end{array}\right),$ (42)
and we similarly decompose the spin-$3$ auxiliary field in traceless spatial components. We thus define
$\chi_{ab0}=\frac{1}{2c^{2}}\alpha\delta_{ab}+v_{ab},\qquad\chi_{abc}=\frac{3}{4c^{2}}\phi_{(a}\delta_{bc)}+S_{abc},$ (43)
@@ Appendix B Non-relativistic limit
while the remaining components are fixed by the constraint $\eta^{\mu\nu}\chi_{\mu\nu\rho}$, for instance $\chi_{a00}=c^{2}\delta^{bc}\chi_{abc}=\phi_{a}$. For the spin-4 field we define
$\Phi_{abc0}=\frac{3}{4c^{2}}\alpha_{(a}\delta_{bc)}+v_{abc},\qquad\Phi_{abcd}=\frac{3}{8c^{4}}\beta\delta_{(cd}\,\delta_{ab)}+\frac{1}{c^{2}}\phi_{(cd}\,\delta_{ab)}+S_{abcd},\qquad$ (44)
@@
where again the remaining components can be obtained using the traceless constraint for $\Phi_{\mu\nu\rho\sigma}$. In two spatial dimensions, any traceless symmetric tensor has only two independent components. Therefore, for every spatial vector or spatial traceless symmetric tensor of the form $Z_{a_{1}\cdots a_{k}}$ with $k=1,2,\dots$, it is possible to construct a complex field $\mathcal{Z}_{[k]}$ out of its independent components as

### v2

@@ Appendix B Proca-like actions for massive particles of any spin
The value of $c_{2}$ fixed by Eq. (47b) is precisely that setting to zero the right-hand side. This allows one to conclude $\partial_{\mu}\partial\cdot B=0$ and, eventually, $\partial\cdot B=0$ up to constant functions that do not affect the propagated local degrees of freedom. This on-shell constraint is crucial because, after the gauge fixing setting the scalar to zero, the spin-$2$ gauge parameter satisfies $\partial\cdot\xi=0$, so that the divergence of the vector potential becomes gauge invariant. On the other hand, the latter is forced to vanish by the equations of motion, so that one can gauge away the remaining portion of the vector potential by using the algebraic gauge transformation generated by $\xi_{\mu}$. After one reaches the condition $B_{\mu}=0$ by combining the equations of motion and a gauge fixing, Eq. (50a) reduces to the Fierz-Pauli equation in Eq. (11).
Appendix C Non-relativistic limit of the Proca-like actions
In this Appendix we detail the Galilean limit of the higher-spin actions of App. B that we discussed in Sect. IV. We restrict the analysis to $D=2+1$ dimensions and we consider explicitly the examples of fields of spin two and four, while commenting on how to extend the limiting procedure to fields of arbitrary spin.
We split the indices as $\mu=\{0,a\}$ and we consider coordinates $x^{\mu}=\{x^{0}\equiv t,x^{a}\}$, so that the Minkowski metric and its inverse are given by $\eta_{\mu\nu}={\rm diag}(-c^{2},1,1)$ and $\eta^{\mu\nu}={\rm diag}(-1/c^{2},1,1)$, respectively.
Spin-2 case:
We begin by considering the Lagrangian (24) for a massive spin-2 field, which correspond to the $s=2$ and $D=3$ instance of (46). The action is invariant under the gauge transformations (25). Using the $\lambda$-symmetry to set to zero the scalar field $S$, we find the Lagrangian
$\begin{split}&\kappa^{-1}{\cal L}=\frac{1}{2}h^{\mu\nu}\left(\Box h_{\mu\nu}-2\partial_{\mu}\partial\cdot h_{\nu}-m^{2}c^{2}h_{\mu\nu}\right)\\ &+\frac{1}{2}B^{\mu}\left(\Box B_{\mu}-\partial_{\mu}\partial\cdot B\right)+\sqrt{2}mc\,h^{\mu\nu}\partial_{\mu}B_{\nu}.\end{split}$ (52)
We can now split $h_{\mu\nu}$ in traceless spatial components as in Eq. (12), while also splitting the Stueckelberg field as $B_{\mu}=(v,S_{a})$. In $2+1$ dimensions we can then combine the two independent components of the previous traceless tensors of rank $k\geq 1$ into the complex fields introduced in Eq. (27). This allows one to rewrite the Lagrangian (52) (up to total derivatives) as
$\kappa^{-1}\mathcal{L}$ $=-\frac{1}{8c^{6}}\dot{\phi}^{2}-\frac{1}{4c^{4}}|\partial\phi|^{2}+\frac{1}{2c^{2}}|\dot{\mathcal{S}}_{[2]}|^{2}+\frac{1}{2c^{2}}\bar{\partial}\mathcal{S}_{[2]}\bar{\partial}\phi+\frac{m}{\sqrt{2}c^{3}}\phi\dot{v}-\frac{3m^{2}}{8c^{2}}\phi^{2}+\frac{m}{2\sqrt{2}c}\phi\partial\mathcal{S}^{*}_{[1]}-\frac{m^{2}c^{2}}{2}|\mathcal{S}_{[2]}|^{2}$ (53)
$+\frac{1}{4c^{2}}\left(|\dot{\mathcal{S}}_{[1]}|^{2}-2\dot{\mathcal{S}}_{[1]}\bar{\partial}v+|\partial v|^{2}\right)-\frac{1}{4}|\partial\mathcal{S}_{[1]}|^{2}+\frac{1}{16}\left(\partial\mathcal{S}^{*}_{[1]}+\bar{\partial}\mathcal{S}_{[1]}\right)^{2}-\frac{mc}{\sqrt{2}}\bar{\partial}\mathcal{S}_{[2]}\mathcal{S}_{[1]}^{*}+\kappa^{-1}\mathcal{L}_{\mathcal{V}_{[1]}}+\rm{c.c.},$
where c.c. denotes the complex conjugate and we defined the complex derivatives $\partial=\partial_{1}+i\partial_{2}$ and ¯$\bar{\partial}=\partial_{1}-i\partial_{2}$. Furthermore, we have isolated the terms that depend on $\mathcal{V}_{[1]}$ in
$\kappa^{-1}\mathcal{L}_{\mathcal{V}_{[1]}}$ $=\frac{m^{2}}{2}|\mathcal{V}_{[1]}|^{2}+\frac{1}{2c^{2}}|\partial\mathcal{V}_{[1]}|^{2}-\frac{1}{c^{2}}\bar{\partial}\mathcal{S}_{[2]}\dot{\mathcal{V}}_{[1]}^{*}+\frac{1}{2c^{4}}\dot{\mathcal{V}}_{[1]}\bar{\partial}\phi-\frac{1}{8c^{2}}\left(\partial\mathcal{V}^{*}_{[1]}+\bar{\partial}\mathcal{V}_{[1]}\right)^{2}\!-\frac{m}{\sqrt{2}c}\left(\mathcal{V}_{[1]}^{*}\partial v+\dot{\mathcal{S}}_{[1]}\mathcal{V}_{[1]}^{*}\right),$ (54)
Before taking the non-relativistic limit, we introduce the complex fields $\mathcal{Q}_{[2]}$, $\mathcal{Q}_{[1]}$ and $\mathcal{U}_{[1]}$ via the field redefinition (28). In terms of the new fields the Lagrangian takes the form
$\kappa^{-1}\mathcal{L}$ $=-mE_{0}|\mathcal{Q}_{[2]}|^{2}-im\mathcal{Q}_{[2]}\dot{\mathcal{Q}}_{[2]}^{*}-\frac{mE_{0}}{2}|\mathcal{Q}_{[1]}|^{2}-\frac{im}{2}\mathcal{Q}_{[1]}\dot{\mathcal{Q}}_{[1]}^{*}+\frac{m^{2}c^{2}}{4}|\mathcal{Q}_{[1]}|^{2}-\frac{mc}{\sqrt{2}}\bar{\partial}\mathcal{Q}_{[2]}\mathcal{Q}_{[1]}^{*}$ (55)
$-\frac{1}{4}|\partial\mathcal{Q}_{[1]}|^{2}+\frac{1}{16}\left(e^{i(mc^{2}-E_{0})t}\partial\mathcal{Q}^{*}_{[1]}+e^{-i(mc^{2}-E_{0})t}\bar{\partial}\mathcal{Q}_{[1]}\right)^{2}+\kappa^{-1}\mathcal{L}_{\mathcal{U}_{[1]}}+\rm{c.c.}+O(c^{-1}).$
Note that the field redefinition (28) introduced some factors $e^{-i(mc^{2}-E_{0})t}$ that will oscillate in the $c\to\infty$ limit. On the other hand, these terms will decouple in the non-relativistic limit, so that we will ignore this issue as in [53]. After introducing the new variables, the terms in Eq. (54) take the form
$\mathcal{L}_{\mathcal{U}_{[1]}}=\frac{\kappa\,m^{2}}{2}|\mathcal{U}_{[1]}|^{2}-\kappa\,\mathcal{U}_{[1]}^{*}\mathcal{J}_{[1]}+O(c^{-2}),$ (56)
where we have defined
$\mathcal{J}_{[1]}$ $=-\frac{im^{2}c}{\sqrt{2}}\mathcal{Q}_{[1]}+im\bar{\partial}\mathcal{Q}_{[2]}$ (57)
$+\frac{m}{\sqrt{2}c}\left(\partial v+iE_{0}\mathcal{Q}_{[1]}+\dot{\mathcal{Q}}_{[1]}\right).$
To proceed, we note that for large values of $c$ the field equations for $\mathcal{U}_{[1]}^{*}$ lead to $\mathcal{U}_{[1]}=\frac{1}{m^{2}}\mathcal{J}_{[1]}+O(c^{-2})$, and therefore integrating out $\mathcal{U}_{[1]}$ yields the following effective Lagrangian
$\mathcal{L}^{eff}_{\mathcal{U}_{[1]}}$ $=-\frac{1}{2}|\partial\mathcal{Q}_{[2]}|^{2}-\frac{m^{2}c^{2}}{4}|\mathcal{Q}_{[1]}|^{2}+\frac{mc}{\sqrt{2}}\mathcal{Q}_{[1]}\partial\mathcal{Q}_{[2]}^{*}$ (58)
$+\frac{im}{2}\mathcal{Q}_{[1]}\left(\bar{\partial}v+\dot{\mathcal{Q}}_{[1]}^{*}-iE_{0}\mathcal{Q}_{[1]}^{*}\right)+O(c^{-1})$
When replacing this expression for $\mathcal{L}_{\mathcal{U}_{[1]}}$ in the Lagrangian (55), the terms multiplied by powers of $c$ that diverge in the limit $c\to\infty$ cancel, and $\mathcal{Q}_{[2]}$ decouples from the other fields. Thus, in the limit $c\to\infty$, up to total derivatives, we find the spin-2 Schroedinger Lagrangian of Eq. (18),
$\mathcal{L}=2i\kappa m\dot{\mathcal{Q}}_{[2]}\mathcal{Q}_{[2]}^{*}-\kappa|\partial\mathcal{Q}_{[2]}|^{2}-2\kappa mE_{0}|\mathcal{Q}_{[2]}|^{2}+\dots,$ (59)
where the dots denote terms that do not depend on $\mathcal{Q}_{[2]}$. Using the fact that up to total derivatives $\epsilon_{bc}Q_{ab}\dot{Q}_{ac}=-2i\dot{\mathcal{Q}}_{[2]}\mathcal{Q}^{*}_{[2]}$, the $\mathcal{Q}_{[2]}$-dependent part of this Lagrangian can be put in the form of Eq. (4), with coefficients $k_{1}=-8\kappa m$, $k_{2}=4\kappa$ and $k_{3}=2\kappa mE_{0}$.
Spin-4 case:
We begin by parametrizing the fields entering the Lagrangian (46) with $s=4$ in terms of traceless spatial components. The full set of relativistic fields can be rewritten in terms of a set of six symmetric and traceless tensors $\{S_{abcd},S_{abc},S_{ab},v_{abc},v_{ab},\phi_{ab}\}$, four vectors $\{S_{a},v_{a},\phi_{a},\alpha_{a}\}$ and five scalar fields $\{\beta,\alpha,v,\phi,S\}$ as follows. For the auxiliary fields of rank one and two we use the same notation as in our analysis of the spin-2 case and we similarly decompose the rank-three auxiliary field in traceless spatial components. We thus define
$\varphi_{ab0}=\frac{1}{2c^{2}}\alpha\delta_{ab}+v_{ab},\quad\varphi_{abc}=\frac{3}{4c^{2}}\phi_{(a}\delta_{bc)}+S_{abc},$ (60)
@@ Appendix C Non-relativistic limit of the Proca-like actions
while the remaining components are fixed by the constraint $\eta^{\mu\nu}\varphi_{\mu\nu\rho}=0$. For the spin-4 field we use the same decomposition as in Eq. (20) with the identification $\varphi_{\mu\nu\rho\sigma}=\Phi_{\mu\nu\rho\sigma}$.
@@
We now profit again from the fact that in two spatial dimensions any traceless symmetric tensor has only two independent components. Therefore, for every spatial traceless symmetric tensor of the form $Z_{a_{1}\cdots a_{k}}$ one can build a complex field $\mathcal{Z}_{[k]}$ out of its independent components as